### 10 Mar 2026 v2

# On the generalized circular projected Cauchy distribution

Abstract
Tsagris and Alzeley, (2025) proposed the generalized circular projected Cauchy distribution, whose special case is the wrapped Cauchy distribution. In this paper we first derive the relationship with the wrapped Cauchy distribution, and then we attempt to characerize the distribution. We provide a non-analytical formula for the mean resultant length, and analytical form for the cumulative probability function and the the entropy. We propose a log-likelihood ratio test for the equality of two angular means, without assuming equality of thew concentration parameters, and briefly mention the addition of circular and simplicial predictors. Simulation studies illustrate a) the performance of the log-likelihood ratio test when one falsely assumes that the true distribution is the wrapped Cauchy distribution, and b) the empirical rate of covergence of the regression coefficients. Using a real data analysis example we correct a mistake in the regression setting.
Keywords: Circular data; mean resultant length, hypothesis test
MSC: 62H11, 62H10
1 Introduction
Directional data refers to multivariate data with a unit norm, whose sample space can be expressed as:
[…]
where $\left|\left|.\right|\right|$ denotes the Euclidean norm. Circular data, when $d=2$, lie on a circle. Circular data are encountered in various disciplines, such as political sciences (Gill and Hangartner,, 2010), criminology (Shirota and Gelfand,, 2017), biology (Landler et al.,, 2018), ecology (Horne et al.,, 2007), and astronomy (Soler et al.,, 2019), to name but a few.
A large class of distributions has been proposed, see Tsagris and Alzeley, (2025) for a short list. In this paper we focus on the generalized circular projected Cauchy distribution (GCPC) proposed by Tsagris and Alzeley, (2025). Specifically we show the relationship with the wrapped Cauchy (WC) distribution and further propose a log-likelihood ratio test for the equality of two independent angular means.
The next section briefly presents the GCPC distribution, derives the relationship with the WC distribution and proposes the log-likelihood ratio test. Then, simulation studies follow and the conclusions that close the paper.
2 The GCPC distribution
[…]
$f({\bf y})=\int_{0}^{\infty}r^{d-1}f(r{\bf y})dr.$ (1)
The probability density function of the bivariate Cauchy distribution, with some location vector $\boldsymbol{\mu}$ and scatter matrix $\boldsymbol{\Sigma}$, is given by
$f({\bf x})=\frac{1}{2\pi|\boldsymbol{\Sigma}|^{1/2}}\left[1+\left({\bf x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left({\bf x}-\boldsymbol{\mu}\right)\right]^{-3/2}.$ (2)
By substituting (2) into (1) and evaluating the integral, Tsagris and Alzeley, (2025) derived the termed the Circular Projected Cauchy (CPC) distribution.
$f({\bf y})$ $=$ $\int_{0}^{\infty}\frac{r}{2\pi|\boldsymbol{\Sigma}|^{1/2}}\left(1+r^{2}{\bf y}^{\top}\boldsymbol{\Sigma}^{-1}{\bf y}-2r{\bf y}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}+\boldsymbol{\mu}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}\right)^{-3/2}dr$ (3)
$=$ $\frac{1}{2\pi|\boldsymbol{\Sigma}|^{1/2}\left(B\sqrt{\Gamma^{2}+1}-A\sqrt{B}\right)},$
where
$A={\bf y}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu},\ B={\bf y}^{\top}\boldsymbol{\Sigma}^{-1}{\bf y},\ \text{and}\ \Gamma^{2}=\boldsymbol{\mu}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\mu}.$
It is important to note that ${\bf y}\in\mathbb{S}^{1}$, while $\bm{\mu}\in\mathbb{R}^{2}$.
(Tsagris and Alzeley,, 2025) relaxed this strict assumption by employing one of the conditions imposed in Paine et al., (2018), that is, $\boldsymbol{\Sigma\mu}=\boldsymbol{\mu}$, but not $|\boldsymbol{\Sigma}|=1$. This condition implies that the one eigenvector $\boldsymbol{\xi}_{2}$ of $\boldsymbol{\Sigma}$ is the normalised location vector $\boldsymbol{\xi}_{2}=\left(\mu_{1},\mu_{2}\right)^{\top}/\gamma$, while the other eigenvector can be defined up to the sign as $\boldsymbol{\xi}_{1}=\left(-\mu_{2},\mu_{1}\right)^{\top}/\gamma$ or $\boldsymbol{\xi}_{1}=\left(\mu_{2},-\mu_{1}\right)^{\top}/\gamma$. The eigenvalue corresponding to the location vector is equal to 1, while the other eigenvalue is equal to $\lambda$, hence $|\boldsymbol{\Sigma}|=\lambda>0$ and the inverse of the scatter matrix is given by
$\boldsymbol{\Sigma}^{-1}=\frac{1}{\gamma^{2}}\left(\begin{array}[]{cc}\mu_{1}^{2}+\mu_{2}^{2}/\lambda&\mu_{1}\mu_{2}\left(1-1/\lambda\right)\\ \mu_{1}\mu_{2}\left(1-1/\lambda\right)&\mu_{2}^{2}+\mu_{1}^{2}/\lambda\end{array}\right)=\boldsymbol{\xi}_{1}\boldsymbol{\xi}_{1}^{\top}/\lambda+\boldsymbol{\xi}_{2}\boldsymbol{\xi}_{2}^{\top}.$ (6)
Thus, (3) becomes
$f({\bf y})=\frac{1}{2\pi\lambda^{1/2}\left(B\sqrt{\gamma^{2}+1}-a\sqrt{B}\right)}.$ (7)
Utilising (6) and after some calculations, the density in (7) may also be expressed in polar coordinates by
$f(\theta)$ $=$ $\frac{\left(2\pi\lambda^{1/2}\right)^{-1}}{\left[\left(\cos^{2}(\theta-\omega)+\frac{\sin^{2}(\theta-\omega)}{\lambda}\right)\sqrt{\gamma^{2}+1}-\gamma\cos(\theta-\omega)\sqrt{\cos^{2}(\theta-\omega)+\frac{\sin^{2}(\theta-\omega)}{\lambda}}\right]}$ (8)
$=$ $\frac{1}{2\pi\lambda^{1/2}\left(b\sqrt{\gamma^{2}+1}-a\sqrt{b}\right)},$
where $a=\gamma\cos\phi,\ b=\cos^{2}\phi+\frac{\sin^{2}\phi}{\lambda},\ \phi=\theta-\omega$, with $\lambda>0$, $\gamma\geq 0$ and $\omega\in[-\pi,\pi]$.
The GCPC distribution exhibits symmetry with respect to $\omega$ such that $f(\theta-\omega)=f(\omega-\theta)$, and is thus symmetrical about $\omega+\pi$. The density of the GCPC may be re-written as
$f(\theta)=\frac{1}{2\pi\sqrt{\lambda}\sqrt{b}D}=\frac{1}{2\pi\sqrt{1+\left(\lambda-1\right)\cos^{2}{\phi}}\cdot D},$ (9)
where
$D=c\sqrt{b}-\gamma\cos\phi\ \text{and}\ c=\sqrt{\gamma^{2}+1}.$
It is important to note that if $\lambda=1$, the GCPC distribution reduces to the circular independent projected Cauchy (CIPC) distribution (Tsagris et al.,, 2025)
$f_{CIPC}({\bf\theta})=\dfrac{1}{2{\pi}\left(\sqrt{\gamma^{2}+1}-\gamma\cos{\left(\theta-\omega\right)}\right)},$ (10)
which is the WC distribution (Mardia and Jupp,, 2000, pg. 51) with a different parameterization
$f_{WC}(\theta)=\frac{1-\delta^{2}}{2\pi\left[1+\delta^{2}-2\delta\cos\left(\theta-\omega\right)\right]},$ (11)
where $\gamma=\frac{2\delta}{1-\delta^{2}}$ or, conversely, $\delta=(\sqrt{\gamma^{2}+1}-1)/\gamma$ (Tsagris and Alzeley,, 2025).
2.1 Relationship with the CIPC distribution
Theorem 2.1.
[…]
$\arctan 2(y,x)=\left\{\begin{array}[]{ll}\arctan(y/x)&\text{if}\ x>0,\\ \arctan(y/x)+\pi&\text{if}\ x<0\ \text{and}\ y\geq 0,\\ \arctan(y/x)-\pi&\text{if}\ x<0\ \text{and}\ y<0,\\ \pi/2&\text{if}\ x=0\ \text{and}\ y>0,\\ -\pi/2&\text{if}\ x=0\ \text{and}\ y<0,\\ \text{undefined}&\text{if}\ x=y=0.\end{array}\right$. (18)
Proof.
The proof is straightfoward by application of the change-of-variables formula to obtain
$f_{\Psi}(\psi)=f_{\Theta}(\theta)\cdot\left|\frac{d\theta}{d\psi}\right|=\frac{1}{2\pi\sqrt{\lambda}\,\sqrt{b}\,D}\cdot\sqrt{\lambda}\cdot b=\frac{1}{2\pi\!\left(\sqrt{\gamma^{2}+1}-\gamma\cos\psi\right)}.$
2.2 Probabilities
Following Tsagris and Alzeley, (2025), instead of the cumulative probability function we derive the probability included within a given interval $(a,b)$, and provide an alternative formula.
$P(a\leq\theta\leq b)=\int_{a}^{b}f_{GCPC}(t)dt=\int_{-\pi}^{b}f_{GCPC}(t)dt-\int_{-\pi}^{a}f_{GCPC}(t)dt=F_{GCPC}(b)-F_{GCPC}(a).$
Applying Theorem 2.1 the formula above becomes:
$P(a\leq\theta\leq b)$ $=$ $\int_{-\pi}^{\psi(b)}f_{WC}(\psi)\,d\psi-\int_{-\pi}^{\psi(a)}f_{WC}(\psi)\,d\psi=F_{WC}(\psi(b))-F_{WC}(\psi(a))$
$=$ $\left[\frac{1}{2}+\frac{1}{\pi}\arctan\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(b)}{2}\right)\right]-\left[\frac{1}{2}+\frac{1}{\pi}\arctan\!\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(a)}{2}\right)\right]$
$=$ $\frac{1}{\pi}\left[\arctan\!\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(b)}{2}\right)-\arctan\left(\frac{1+\delta}{1-\delta}\tan\frac{\psi(a)}{2}\right)\right]$
where $\psi(t)=\arctan 2\left(\sin(t-\omega),\,\sqrt{\lambda}\cos(t-\omega)\right)$.
2.3 Mean resultant length
The mean resultant length is defined as $\rho=E\left[\cos{\left(\theta-\omega\right)}\right]$. Based on this, one may compute the circular variance as $1-\rho$, and the circular standard deviation as $\left(-2\log{\rho}\right)^{1/2}$. Using Theorem 2.1 we may express this expectation as
$\rho=\frac{1-\delta^{2}}{\pi}\int_{0}^{\pi}\frac{\cos\psi}{\sqrt{\cos^{2}\psi+\lambda\sin^{2}\psi}}\cdot\frac{1}{1+\delta^{2}-2\delta\cos\psi}\,d\psi.$
After some algebra we end up with the following representation
$\rho=\frac{2(1+\delta)}{\pi(1-\delta)\sqrt{\lambda}}\;\Pi\!\left(\left(\frac{1+\delta}{1-\delta}\right)^{2}\frac{1}{\lambda}-1\;\Bigg|\;1-\frac{1}{\lambda^{2}}\right),$
where $\Pi(a|b)$ is the complete elliptic integral of the third kind (Ward,, 1960). We still see that there is no closed form solution for $\rho$, unless $\lambda=1$, and this applies to higher trigonometric moments. Figure 1(a) visualizes the values of $\rho$ for a grid of values of $\gamma$ and $\lambda$. We observe that $\rho$ increases with increasing $\gamma$ and decreases with increasing $\lambda$ values.
2.4 Entropy
Theorem 2.2.
The entropy of the GCPC distribution is given by
$H_{GCPC}(\omega,\gamma,\lambda)=\log{\left\{\frac{8\pi\sqrt{\lambda}\,(1-\delta^{2})}{\bigl[\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}\bigr]^{2}}\right\}},$
where $\delta$ was defined earlier.
Proof.
The entropy is defined as $H=-\int_{-\pi}^{\pi}f(\theta)\log f(\theta)\,d\theta$. Taking the logarithm of (9), we have: $\log f_{GCPC}=-\log(2\pi)-\frac{1}{2}\log\lambda-\frac{1}{2}\log b-\log D$. Using the change of variables from Theorem 2.1, we have $f_{GCPC}(\theta)\,d\theta=f_{WC}(\psi)\,d\psi$ and
$\log D=\frac{1}{2}\log b+\log(c-\gamma\cos\psi)\ \ \text{and}\ \ b\left[\phi(\psi)\right]=\frac{1}{1+(\lambda-1)\sin^{2}\psi},$
where $c-\gamma\cos\psi=(1+\delta^{2}-2\delta\cos\psi)/(1-\delta^{2})$.
By changing variables and substituting the result into the entropy we obtain:
$H_{GCPC}=\log(2\pi)+\frac{1}{2}\log\lambda+E_{WC}\left(\log b\right)+\int_{-\pi}^{\pi}f_{WC}(\psi)\log(c-\gamma\cos\psi)\,d\psi.$
The last integral equals $H_{WC}-\log{\left(2\pi\right)}$, thus
$H_{GCPC}=H_{WC}+\frac{1}{2}\log\lambda-E_{WC}\left[\log(1+(\lambda-1)\sin^{2}\psi)\right].$ (19)
We know that the entropy of the WC distribution is $H_{WC}=\log{\left[2\pi\left(1-\delta^{2}\right)\right]}$. Let us now write $1+(\lambda-1)\sin^{2}\psi=\frac{1+\lambda}{2}(1-\alpha\cos 2\psi)$ with $\alpha=(\lambda-1)/(\lambda+1)$. Then, $1-\alpha\cos 2\psi=(1+r_{1}^{2}-2r_{1}\cos 2\psi)/(1+r_{1}^{2})$ where
$r_{1}=\frac{1-\sqrt{1-\alpha^{2}}}{\alpha}=\frac{\sqrt{\lambda}-1}{\sqrt{\lambda}+1}.$
Using the Fourier identity $\log(1+\delta^{2}-2\delta\cos\psi)=-2\sum_{n=1}^{\infty}\frac{\delta^{n}}{n}\cos(n\psi)$ we get $E_{\mathrm{WC}}[\cos(n\psi)]=\delta^{n}$. Applying the same Fourier identity at frequency $2n$, with $E_{WC}\left[\cos(2n\psi)\right]=\delta^{2n}$:
$E_{WC}\left[\log(1-\alpha\cos 2\psi)\right]=-\log(1+r_{1}^{2})+2\log(1-r_{1}\delta^{2}).$
Using $\log\frac{1+\lambda}{2}-\log(1+r_{1}^{2})=2\log\frac{\sqrt{\lambda}+1}{2}$ and $1-r_{1}\delta^{2}=\frac{\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}}{\sqrt{\lambda}+1}$ we obtain
$E_{WC}\left[\log(1+(\lambda-1)\sin^{2}\psi)\right]=2\log\frac{\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}}{2}.$
Substituting the above result into (19) we obtain
$H_{GCPC}(\omega,\gamma,\lambda)=\log{\left\{\frac{8\pi\sqrt{\lambda}\,(1-\delta^{2})}{\bigl[\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}\bigr]^{2}}\right\}}.$
Note that if $\lambda=1$: $H_{GCPC}=\log(2\pi(1-\delta^{2}))=H_{WC}$, and if $\delta=0$: $H_{GCPC}=\log(8\pi\sqrt{\lambda}/(\sqrt{\lambda}+1)^{2})\leq\log(2\pi)$. Figure 1(b) visualizes the values of $H$ for a grid of values of $\gamma$ and $\lambda$. We observe that, in cotrast to $\rho$, $H$ increases with increasing $\lambda$ and decreases with increasing $\gamma$ values.
Refer to caption Refer to caption
(a) $\rho$ (b) $H$
Figure 1: Values of the mean resultant length, $\rho$ and entropy, $H$, as a function of $\gamma$ and $\lambda$.
2.5 Kullback-Leibler divergence between GCPC and CIPC
[…]
$KL(GCPC\|CIPC)$ $=$ $-\frac{1}{2}\log\lambda+2\log\!\left(\frac{\sqrt{\lambda}(1-\delta^{2})+1+\delta^{2}}{2}\right)-\log(1-\delta^{2})$
$+E_{WC}\!\left[\log\!\left(\sqrt{\gamma^{2}+1}-\frac{\gamma\cos\psi}{\sqrt{1+(\lambda-1)\sin^{2}\psi}}\right)\right],$
where the expectation is taken with respect to $\psi\sim WC(0,\delta)$.
2.6 Hypothesis testing and confidence intervals
2.6.1 Hypothesis test for the location parameter of one sample
In order to test whether the location parameter equals some pre-specified value we will employ the log-likelihood ratio test. Under the $H_{0}$, $\omega=\omega_{0}$, and so we need to maximize the restricted log-likelihood, $\ell_{0}$, with respect to the other two parameters, $\gamma$ and $\lambda$, and estimate their values, $\tilde{gamma}$ and $\tilde{\lambda}$. Under the $H_{1}$, $\omega\neq\omega_{0}$, and we maximize the unrestricted log-likelihood, $\ell_{1}$, to obtain $\hat{\omega}$, $\hat{\gamma}$ and $\hat{\lambda}$. Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\omega},\hat{\lambda}\right)-\ell_{0}\left(\omega_{0},\tilde{\gamma},\tilde{\lambda}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
2.6.2 Confidence interval for a location parameter
Using the log-likelihood ratio test, we can construct asymptotic confidence intervals for the true location parameter $\omega$ of a sample by searching for the pair of values for which it holds that (Cox and Hinkley,, 1979): $\ell\left(\omega,\hat{\gamma},\hat{\lambda}\right)>\ell\left(\hat{\omega},\hat{\gamma},\hat{\lambda}\right)-\frac{\chi^{2}_{1}}{2}$.
2.6.3 Equality of two location parameters
To test the equality of two location parameters, without assuming equality of the parameters $\gamma$ and $\lambda$, we will perform a log-likelihood ratio test [^1] The function to perform the log-likelihood ratio test is available in the R package Directional (Tsagris et al.,, 2026). following (Tsagris et al.,, 2025). Assume we have circular observations from two independent samples, $(\theta_{11},\ldots,\theta_{1{n_{1}}})$ and $(\theta_{11},\ldots,\theta_{1{n_{2}}})$, where $n_{1}$ and $n_{2}$ denote the sample sizes of the two samples.
Under $H_{0}$, $\omega_{1}=\omega_{2}=\tilde{\omega}$, and by using Eq. (8), the log-likelihood is written as
$\ell_{0}\left(\tilde{\omega},\tilde{\lambda}_{1},\tilde{\lambda}_{2},\tilde{\gamma}_{1},\tilde{\gamma}_{2}\right)$ $=$ $-(n_{1}+n_{2})\log(2\pi)-\frac{n_{1}}{2}\log(\tilde{\lambda}_{1})-\frac{n_{2}}{2}\log(\tilde{\lambda}_{2})$
$-\sum_{i=1}^{n_{1}}\log{\left(\tilde{b}_{1i}\sqrt{\tilde{\gamma}_{1}^{2}+1}-\tilde{a}_{1i}\sqrt{\tilde{b}_{1i}}\right)}-\sum_{i=1}^{n_{2}}\log{\left(\tilde{b}_{2i}\sqrt{\tilde{\gamma}_{2}^{2}+1}-\tilde{a}_{2i}\sqrt{\tilde{b}_{2i}}\right)},$
where $\tilde{a}_{ji}=e^{\tilde{\gamma}_{j}}\cos{\left(\theta_{ji}-\tilde{\omega}\right)}$ and $\tilde{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\tilde{\omega}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\tilde{\omega}\right)}{\tilde{\lambda}_{j}}$, for $j=1,2$.
Under $H_{1}$, $\omega_{1}\neq\omega_{2}$, and the log-likelihood is written as
$\ell_{1}\left(\hat{\omega},\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\gamma}_{1},\hat{\gamma}_{2}\right)$ $=$ $-(n_{1}+n_{2})\log(2\pi)-\frac{n_{1}}{2}\log(\hat{\lambda}_{1})-\frac{n_{2}}{2}\log(\hat{\lambda}_{2})$
$-\sum_{i=1}^{n_{1}}\log{\left(\hat{b}_{1i}\sqrt{\gamma_{1}^{2}+1}-\hat{a}_{1i}\sqrt{\hat{b}_{1i}}\right)}-\sum_{i=1}^{n_{2}}\log{\left(\hat{b}_{2i}\sqrt{\hat{\gamma}_{2}^{2}+1}-\hat{a}_{2i}\sqrt{\hat{b}_{2i}}\right)},$
where $\hat{a}_{ji}=e^{\hat{\gamma}_{j}}\cos{\left(\theta_{ji}-\hat{\omega}_{j}\right)}$ and $\hat{b}_{ji}=\cos^{2}{\left(\theta_{ji}-\hat{\omega}_{j}\right)}+\frac{\sin^{2}\left(\theta_{ji}-\hat{\omega}_{j}\right)}{\hat{\lambda}_{j}}$, for $j=1,2$.
Under regularity conditions, $\Lambda=2\left[\ell_{1}\left(\hat{\omega},\hat{\lambda}_{1},\hat{\lambda}_{2},\hat{\gamma}_{1},\hat{\gamma}_{2}\right)-\ell_{0}\left(\tilde{\omega},\tilde{\lambda}_{1},\tilde{\lambda}_{2},\tilde{\gamma}_{1},\tilde{\gamma}_{2}\right)\right]\mathrel{\dot{\sim}}\chi^{2}_{1}$.
2.7 Regression modelling revisited
Tsagris and Alzeley, (2025) defined the GCPC regression by considering the bivariate representation (Euclidean coordinates) of the circular data as opposed to their univariate nature (Kato et al.,, 2008). This is akin to the approach employed in the SPML model, as detailed by Presnell et al., (1998) and allows for varying concentration parameter ($\lambda$). The log-likelihood of the GCPC regression model is written as
$\ell_{GCPC}$ $=$ $-\sum_{i=1}^{n}\log{\left(\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}\sqrt{\bm{\mu}_{i}^{\top}\bm{\mu}_{i}+1}-\bm{y}_{i}^{\top}\bm{\mu}_{i}\sqrt{{\bf y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}}\right)}$ (20)
$-\frac{n}{2}\log(\lambda)-n\log{(2\pi)},$
where
$\bm{y}_{i}^{\top}\bm{\Sigma}^{-1}_{i}\left(\bm{B}\right)\bm{y}_{i}$ $=$ $\xi_{1i}\left(\bm{B}\right)\xi_{1i}\left(\bm{B}\right)^{\top}/\lambda+\xi_{2i}\left(\bm{B}\right)\xi_{2i}\left(\bm{B}\right)^{\top}$
$=$ $y_{1i}^{2}\left(\xi_{1i}^{2}/\lambda+\xi_{2i}^{2}\right)+y_{2i}^{2}\left(\xi_{2i}^{2}/\lambda+\xi_{1i}^{2}\right)+2y_{1i}y_{2i}\xi_{1i}\xi_{2i}\left(1-1/\lambda\right)$
and
$\xi_{ji}=\frac{\bm{\beta}_{j}^{\top}x_{i}}{\|\bm{\beta}_{j}^{\top}x_{i}\|}=\frac{\mu_{ji}}{\|\bm{\mu}_{i}\|}=\frac{\mu_{ji}}{\gamma_{i}}\ \ \text{for}\ \ j=1,2,$ (21)
where $\bm{x}_{i}$ denoting the $i$-th observation of the covariate vector.
To maximize the log-likelihood of the GCPC regression model (20) suggested to use multiple starting points, but we propose a faster option. We begin with initial regression coefficients produced by the SPML regression model of Presnell et al., (1998) and then estimate the $\lambda$ parameter using Brent’s algorithm [^2] This is available in R via the buit-in optimize() function. (Brent,, 2013). Then use these as starting points in R’s built-in optim() function.
Tsagris and Alzeley, (2025) did not consider the case of circular-circular regreression, where a covariate $X$ is circular as well. But, this case is easily adopted in the above case scenario by transforming the circular covariate to Euclidean by using $\bm{z}_{i}=\left(\cos{x_{i}},\sin{x_{i}}\right)^{\top}$. In case of multiple circular covariates, all of them are transformed into their Euclidean coordinates and the same link function (21) is used. Simplicial predictors (compositional data) are straighforward to add by using two options. The first is to apply the additive log-ratio transformation (Aitchison,, 2003), so that the regression coefficients sum to zero and have a nice (derivative based) interpretation. In case of zero values present the logarithmic transformation breaks down, so the alternative approach is to transform them via the $\alpha$–transformation (Tsagris et al.,, 2011). The second approach bypasses the problem of zero values but lacks interpretation of the estimated regression coefficients and requires to choose the value of the $\alpha$ parameter. The choice of $\alpha$ may be overcome by choosing a small value of $\alpha$ (say 0.01) to resemble the isometric log-ratio transformation (Egozcue et al.,, 2003).
An extra extension would be to link the $\lambda$ parameter to the covariates as well, but that would increase the computational cost. Another evident approach to make the model non-linear would be to use splines in the covariates as well.
3 Simulation studies
We performed simulation studies to examine the performance of the two sample means test and also the effect of the $\lambda$ parameter on the performance of the log-likelihood test.
3.1 One location parameter
We selected six sample sizes, and generated values from the $GCPC\left(2,3,2\right)$. For each sample size we computed the type I error (assuming $\alpha=0.05$) based on 1,000 replicates. Table 1 presents the results, where we can see that the test slightly overestimates the nominal type I error, but not significantly.
Table 1: Estimated type I error of the log-likelihood ratio test for one location parameter assuming the GCPC distribution.
Sample size 30 50 70 100 150 200
Type I error 0.065 0.059 0.066 0.062 0.067 0.063
3.2 Two location parameters
We compared the test when both samples come from the GCPC distribution and when assuming that they come from the CIPC distribution ($\lambda=1$). Six pairs of unequal sample sizes were used, where the true angular mean was $\omega=2$, but the $\gamma$ and $\lambda$ parameters differed. The parameters for the smaller sample were equal to $\gamma=4$ and $\lambda=1$, while for the larger sample they were equal to $\gamma=2$ and $\lambda=3$. Circular data were generated from two GCPC distributions with the specified parameters and the two log-likelihood ratio tests were performed. This process was repeated 1,000 times and the proportion of rejection of the $H_{0}$ (at the $\alpha=0.05$ level) served as an estimate of the type I error. Table 2 presents the results. The GCPC based log-likelihood ratio always attained the correct size, whereas the CIPC based log-likelihood ratio test overestimated the size of the test.
Table 2: Estimated type I error of the log-likelihood ratio test for the equality of two location parameters, assuming the GCPC distribution and assuming the CIPC distribution.
Sample size GCPC CIPC
(30, 50) 0.054 0.098
(30, 70) 0.056 0.094
(30, 100) 0.060 0.102
(50, 70) 0.048 0.100
(50, 100) 0.059 0.101
(70, 100) 0.061 0.092
3.3 Empirical asymptotics for the regression model
We estimated empirically the order of convergence of the regression coefficients, with a) one circular predictor, b) one continuous predictor and c) the combination of both. The grid for the sample sizes considered ranged from 100 up to 5,000 at an increasing step of 100. For every sample size $n$, data were generated with some pre-specified coefficients and the regression model was fitted.
For the circular predictor, the model was
$\left(y_{1},\ y_{2}\right)=\left(1,\ \cos{u},\ \sin{u}\right)\left(\begin{array}[]{cc}\beta_{01}&\beta_{02}\\ \beta_{11}&\beta_{12}\\ \beta_{21}&\beta_{22}\end{array}\right),$
where $\bm{u}$ was generated from a von Mises distribution with mean equal to 2 and concentration parameter equal to 5. For the continuous predictor, the model was
$\left(y_{1},\ y_{2}\right)=\left(1,\ x\right)\left(\begin{array}[]{cc}\beta_{01}&\beta_{02}\\ \beta_{11}&\beta_{12}\\ \end{array}\right),$
where $\bm{x}$ was generated from an exponential distribution with mean equal to 0.5. Table 3 shows the matrix with the ground truth regression coefficients $\bm{B}$ used. For all cases, the $\lambda$ values of the GCPC distribution used to generate the circular responses were equal to 0.5 and 5.
The discrepancy between the observed and the estimated regression coefficients was measured using the squared Frobenius norm. This process was repeated 50 times and the average squared Frobenius norm ($\bar{F}^{2}$) was stored. Then, we estimated the coefficient of the following regression model:
$\log{\bar{F}^{2}}\sim a+b\log{n}.$
The estimated $b$ served as the empirical rate of convergence of the regression coefficients.
Table 3 contains the matrix $\bm{B}$ of the regression parameters used in the simulation study. Figure 2 visualizes the rates when $\lambda=0.5$ and $\lambda=5$. For all cases the estimated slope is very close to 1, indicating that the convergence rate of the estimated regression coefficients is $n^{-1}$.
Table 3: Matrix $\bm{B}$ of the regression coefficients used in the simulations. The first two columns correspond to the coefficients of the model with the circular predictor, the second two to the model with the continuous predictor and the last two to the model with both of them. In the last case $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ refer to the circular predictor and $\bm{\beta}_{3}$ to the continous predictor.
Predictor Circular Continuous Circular and continuous
$\cos(y)$ $\sin(y)$ $\cos(y)$ $\sin(y)$ $\cos(y)$ $\sin(y)$
$\bm{\beta}_{0}$ 1.874 2.550 1.641 1.949 1.859 3.508
$\bm{\beta}_{1}$ -1.473 -2.023 -0.101 -0.119 -1.500 -2.914
$\bm{\beta}_{2}$ -1.157 -1.554 -0.960 -1.855
$\bm{\beta}_{3}$ -0.246 -0.181
Circular covariate
Refer to caption Refer to caption
(a) $\lambda=0.5$ (b) $\lambda=5$
Continuous covariate
Refer to caption Refer to caption
(c) $\lambda=0.5$ (d) $\lambda=5$
Circular and continuous covariate
Refer to caption Refer to caption
(e) $\lambda=0.5$ (f) $\lambda=5$
Figure 2: Average squared Frobenius norm $\bar{F}^{2}$ versus sample size, both in log-scale for the circular, the continuous and the combination of both.
4 Real data analysis
We used the same dataset used in Tsagris and Alzeley, (2025), the speed.wind2 data that is available in the R package NPCirc (Oliveira et al.,, 2014). This dataset consists of 199 hourly observations of wind direction and wind speed in winter season (from November to February) from 2003 until 2012 in the Atlantic coast of Galicia (Spain).
4.1 Maximum likelihood estimation
Table 4 presents the estimated parameters of the CIPC and GCPC models. Application of the log-likelihood ratio test to discriminate between the two models (Tsagris and Alzeley,, 2025) clearly favors the GCPC model. Figure 3(a) visualizes the data, it is the circular plot with the mean directions showed with arrows. Evidently the data look bimodal and the CIPC has chosen one of the two modes, whereas the GCPC has chosen a direction that ”cuts” the data into two. The density plot on Figure 3(b) shows that the GCPC density has a similar pattern to the one produced by the kernel density estimate, in contrast to the CIPC distribution that has failed to fit the data. This is more evident in the circular density plot (Figure 3(c)). The CIPC distribution has a circular shape, whereas the GCPC distribution has an elliptical shape, capturing the shape of the data more accurately. Finally, Figure 3(d) visualizes the 95% confidence interval for the mean direction.
Table 4: Estimated parameters of the CIPC and GCPC models applied to the wind direction data.
Model $\hat{\omega}$ $\hat{\gamma}$ $\hat{\lambda}$ Log-likelihood
CIPC 0.603 0.203 -363.930
GCPC 5.587 0.050 4.21 -337.739
Refer to caption Refer to caption
(a) Circular plot (b) Density plot
Refer to caption Refer to caption
(c) Circular density plot (d) 95% confidence interval for the mean direction
Figure 3: (a) Circular plot of the speed direction data. (b) Density plot of the kernel density estimate, and the fitted CIPC and GCPC distributions. (c) Circular density plot of the two fitted distributions. The green line is the estimated mean direction based on the CIPC, whereas the blue lines is the estimated mean direction based on the GCPC model.
4.2 Regression
We spotted a mistake in the regression modelling of Tsagris and Alzeley, (2025), where the continuous predictor is the speed of the wind. Table 5 shows the regression coefficients computed in Tsagris and Alzeley, (2025) and the correct estimates. Note that this time the sign of the slope coefficients matches that of the CIPC and of the spherically projected multivariate linear regression models models. The $\lambda$ coefficient still remains statistically significantly far from 1, and the log-likelihood ratio test still favors the GCPC regression over the CIPC regression model.
Table 5: Estimated regression parameters (their standard errors appear within parentheses) for the GCPC regression model fitted to the wind data. On the right is the corrected version of the GCPC regression model.
Model GCPC (Tsagris and Alzeley,, 2025) GCPC
$\cos(y)$ $\sin(y)$ $\cos(y)$ $\sin(y)$
$\hat{\bm{\alpha}} -$0.042 (0.0013) -0.045 (0.0023) 0.164 (0.0012) 0.195 (0.0012)
$\hat{\bm{\beta}}$ 0.009 (0.0005) 0.010 (0.0009) -0.010 (0.0022) -0.012 (0.0021)
Loglik -331.846 -335.219
$\hat{\rho}$ 0.203 (0.039) 0.226 (0.055)
5 Conclusions
We derived the relationship between the GCPC and the CIPC (reparameterized wrapped Cauchy) distribution, provided a not so convenient formula for the mean resultant length, an alternative analytical formula to compute probabilities, and an analytical formula for the entropy of the GCPC. We further proposed two log-likelihood ratio tests for hypothesis testing with one or two location parameters. For the two location parameters testing case, the test is size correct when the data are wrongfully assumed to follow the CIPC distribution. Finally we corrected a mistake in the regression setting and proposed a faster altenative than the one proposed by Tsagris and Alzeley, (2025).